# Modelling the evolution and influence of dust in cosmological simulations that include the cold phase of the interstellar medium

### v1

@@ 4.5 Dust-to-gas ratio
Figure 9: The dust-to-gas ($\mathcal{DTG}$) ratio of relatively active (sSFR$={\rm SFR}/M_{\star}>0.01\;{\rm Gyr}^{-1}$) galaxies as a function of their gas-phase oxygen abundances, measured in cool ($\log_{10}T/{\rm K}<4.5$), dense ($\log_{10}n_{\rm H}/{\rm cm^{-3}}>-1$) gas, where ${\rm O/H}$ values are aggregated as the total linear ratio of oxygen to hydrogen in selected, in-aperture gas. These are compared to a compilation of observations (Konstantopoulou et al., 2023; Cigan et al., 2021; De Vis et al., 2019; Galametz et al., 2011; Galliano et al., 2008). Blue lines show median values while blue data points show individual galaxies. The solid line shows the median for galaxies with $M_{\star}>10^{9}~{}{\rm M_{\odot}}$ and ±$\pm 1\sigma$ ranges indicated by the shaded region and blue squares showing individual galaxies. The dotted line shows the median relation for all processed galaxies, with blue dots indicating galaxies outside the high mass selection. We also use (KDE-smoothed) contours (with 0.46 dex spacing) to indicate the histogram of De Vis et al. (2019) data using an alternative metallicity calibration (T04, Tremonti et al. 2004), as opposed to their fiducial one (PG16R, Pilyugin & Grebel 2016). Metallicity calibrations are investigated further in Appendix C.
@@
For $\mathcal{DTG}$, plotting values as a function of gas-phase metallicity can reveal how dust content varies with the ongoing enrichment of the ISM. Observationally, this metallicity is typically measured in the diffuse gas phase via oxygen emission lines, taken as a proxy for total metallicity, $12+\log_{10}({\rm O/H})$. For our simulated galaxies, we compute this per subhalo for our simulated galaxies in 50 pkpc apertures about the most-bound particles, measuring the diffuse oxygen fractions in relatively cold, dense gas ($\log_{10}T/{\rm K}<4.5$, $\log_{10}n_{\rm H}/{\rm cm^{-3}}>-1$) gas, to approximate the regions probed observationally. Here, we limit the galaxy selection to actively star-forming galaxies (with specific star formation rates ${\rm SFR}/M_{\star}>0.01\;{\rm Gyr}^{-1}$), with $12+\log_{10}({\rm O/H})$ abundances aggregated linearly over the subhalo material (such that O/H is the ratio of total oxygen to hydrogen across all the cool-dense, in-aperture gas). Another important consideration for the comparison in this case is that with $12+\log_{10}({\rm O/H})$ as the independent variable, there can be considerable blending in particle resolution across the $x$-axis, particularly due to low-mass galaxies scattered to high metallicity bins. By default, we apply a stellar mass cut of $\log_{10}(M_{\star}/{\rm M_{\odot}})>9$ (corresponding to galaxies resolved by $\gtrapprox 500$ star particles) to our primary median relation (solid line), while also plotting a median relation using a lower $\log_{10}(M_{\star}/{\rm M_{\odot}})>8$ cut ($\gtrapprox 50$ star particles) to show the influence of less well resolved systems.
@@
Comparing the relation for better resolved ($\gtrapprox 500$ star particles) galaxies with a broader selection using a lower resolution threshold ($\gtrapprox 50$ star particles), we see that for ($8.412+\log_{10}({\rm O/H})>$ 8.4) the broader selection is marginally lower ($\lesssim 0.1$ dex) owing to the inclusion of less-resolved galaxies. We begin to see the broad selection $\mathcal{DTG}$s diverge to lower values at $12+\log_{10}({\rm O/H})\sim 8.25$, falling super-linearly to $12+\log_{10}({\rm O/H})\lesssim 8.1$, and approximately linearly again below that. We note that this transition feature appears to be resolution-dependent, shifting to lower metallicities at higher resolution (Vijayan et al. in prep.). For the narrower selection $12+\log_{10}({\rm O/H})\lesssim 8.2$ we run out of galaxies, owing to a lack of low-metallicity massive objects.
@@ 4.6 Dust grain sizes
Another property we can probe with our dust evolution model is the grain size distribution. In Fig 10, we plot the logarithmic mass ratio of small grains relative to large grains for galaxies, $\log_{10}S/L$. We first focus on ratios computed for all gas within the 50 pkpc exclusive aperture defining each galaxy (blue lines, shading and points). This provides a metric for the grain size distribution in galaxies, where positive values are small-grain mass dominated and negative values are large-grain mass dominated. We see that for our Fiducial model, median galaxy dust masses are dominated by large (0.1 µm) grains for $\log_{10}(M_{\star}/{\rm M_{\odot}})>8.5$. In general, we see a weakly decreasing trend for our median values, such that the average galaxy goes from rough parity between the mass in small and large grains, to most massive galaxies having $\approx 60\%$ of their dust mass in large grains. This trend is significantly smaller than the scatter, however, which is of order 0.4-0.6 dex, and exhibiting diverse extremes; extending between close to an order of magnitude more mass in large grains, to around $50\%$ more mass in small grains. This diversity in the size distribution is profound, and will necessarily impact the extinction properties of grains (which are particularly relevant for post-processing observables from simulated galaxies), as well as the rates of dust-related cooling and heating processes at a fixed $\mathcal{DTG}$ ratio.
@@
To investigate this apparent inconsistency further, we can look at the small-to-large grain ratio within a more stringent gas selection. With $S/L$ ratios showing strong trends in gas density and phase (see e.g. Fig. 5), we over-plot medians for H2-weighted average $S/L$ values (red), as well a after imposing lower limits on the hydrogen number density, $n_{\rm H}>10\;{\rm cm}^{-3}$ and $n_{\rm H}>100\;{\rm cm}^{-3}$. We observe that considering H2-weighted gas does generally push $S/L$ to lower values, but only by $\approx 0.1$ dex for $\log_{10}(M_{\star}/{\rm M_{\odot}})>9$. Considering high-density gas selections has a stronger impact, with $n_{\rm H}>10\;{\rm cm}^{-3}$ gas 0.3-0.5 dex lower than the all-gas ratios, and $n_{\rm H}>100\;{\rm cm}^{-3}$ with an offset of $>1$ dex below all-gas, consistent with the lower sequence of observed values. Along with the gas selection, we note that again resolution appears to play a role here, with the higher m5 (lower m7) resolution yields higher (lower) $S/L$ ratios respectively (Vijayan et al. in prep.).
@@
4.7 Resolved properties
[…]
Finally, we briefly present a gallery of randomly oriented (i.e. box-projected) dust maps for a handful of galaxies taken from our Fiducial model in Fig. 11. These are the 6 highest stellar mass galaxies, defined within a 50 ckpc spherical aperture, that also have $>10\%$ of their baryonic mass constituted by gas. For each galaxy, we map the highest dust surface density ($\Sigma_{\rm d}$) pixels with colour shading, using a kernel density projection of the SPH gas particles, via the swiftgalaxy161616github.com/SWIFTSIM/swiftgalaxy tool. These are plotted alongside neutral and molecular hydrogen contours to show the relatively cold-dense gas regions. Colour shadings illustrate various properties of the dust population; the local $\mathcal{DTZ}$ ratio, the small-to-large grain ratio, the ratio of silicate to graphite/carbon grains and the ratio of forsterite (Mg-endmember) to fayalite (Fe-endmember) silicates.
@@ 5 Discussion & Conclusions
We have presented a model for the life-cycle of dust and its interaction with the physics of the ISM, implemented using the SWIFT SPH code (Schaller et al., 2024), as part of a network of physics modules constituting the COLIBRE suite of galaxy formation simulations (Schaye et al. in prep.). This model uses an established two-size (0.01 ${\rm\upmu m}$ and 0.1 ${\rm\upmu m}$) grain population paradigm (Hirashita, 2015), as well as three chemical species of dust (carbonaceous grain, as well as Mg- and Fe-endmember silicate species), providing a model lightweight enough to be carried particle-by-particle in large SPH cosmological volume simulations. This dust is seeded from stellar channels, budgeting for consistency with overall metal enrichment, and can accrue mass from the diffuse ISM (accretion), change its size distribution (shattering and coagulation) and ultimately be destroyed (sputtering, direct shock annihilation or astration). During this life-cycle, dust properties are coupled to the cooling and heating processes, accounting for the role of dust in e.g. absorbing and re-emitting ionising radiation, providing surfaces for molecule formation and depleting metals from the diffuse phase that would otherwise contribute to cooling and heating processes. We presented results from (25 cMpc)3 moderate resolution ($m_{\rm gas}\approx m_{\rm DM}\approx 10^{6}\;{\rm M_{\odot}}$) simulations showing the effect of parameter and modelling choices on the distribution, depletion and evolution of dust, as well as results comparing galaxy dust distributions and scaling relations at redshift $z=0$ to observations.
Generally, we find that our model produces agreement with a variety of observed properties. Our Fiducial run exhibits a cosmic dust density ($\rho_{\rm c,dust}$) evolution in good agreement with a wide range of observations across different redshift windows and observational probes (Fig. 6). We put particular stock in the recent Chiang et al. (2025) results, which integrate the cosmic infrared background to probe all dust emission, so well-matched to our volume-complete $\rho_{\rm c,dust}$ calculation. This shows a new level of agreement across the redshift range for cosmological models of dust evolution, compared to various literature models we consider. While it is arguable that the predicted $\rho_{\rm c,dust}$ evolution is flatter than observed over the range $20<z<$ 2, on the high side compared to low redshift observations and vice-versa, this is consistent with volume effects in a relatively small 253 cMpc3 box that lacks the massive halos that assemble vigorously at high redshift but are largely evacuated of ISM at low redshift.
@@ Acknowledgements
We gratefully acknowledge everyone who has contributed to the COLIBRE project for their input. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. JT acknowledges support of his STFC Early Stage Research and Development grant (ST/X004651/1). We are grateful to Maarten Baes, Andrea Gebek and Nick Andreadis for helpful discussions in the latter phase of this work. This work made generous use of the pipeline tools developed for the COLIBRE project, representing many years of work; including the swiftsimio (Borrow & Borrisov, 2020) and SOAP181818github.com/openjournals/joss-reviews/issues/7851 tools. SP acknowledges support from the Austrian Science Fund (FWF) through project V 982-N. ABL acknowledges support by the Italian Ministry for Universities (MUR) program ‘Dipartimenti Di Eccellenza 2023-2027’ within the Centro Bicocca dinCosmologia Quantitativa (BiCoQ), and support by UNIMIB’s Fondo Di Ateneo Quota Competitiva (project 2024-ATEQC-0050). CSF acknowledges support from ERC Advanced Investigator Grant, DMIDAS (GA 786910). This project has received funding from the Netherlands Organization for Scientific Research (NWO) through research programme Athena 184.034.002.
@@
Appendix A Extinction Efficiencies and Dust Coupling
@@
Appendix B Additional Parameter Variations
B.1 Smaller small grain sizes
We see that smaller grains shift the depletion-limited $\mathcal{DTZ}$ regime to lower densities, and boost the $S/L$ grain ratio by $\approx 0.1$ dex relative to Fiducial in $-2\lesssim\log_{10}n_{\rm H}/{\rm cm^{-3}}\lesssim 2$. While the change in these $\mathcal{DTZ}$-density relations is relatively small, we see that this acts in concert with the smaller grain size to produce a more marked shift in the H2 transition (vertical line mark) to lower densities; a higher number of grains with higher surface-to-volume ratios at intermediate densities speeds up H2 nucleation.
B.2 Turbulent diffusion
Another process that can influence the evolution of the dust, and its associated influence on gas physics, is turbulent diffusion. As described in 2.6, the diffusion of dust follows that of gas-phase elements described in Correa et al. (2025). The amount of diffusion is modulated by a diffusion constant, $C_{\rm d}$, where elements or grains are diffused more readily for higher values. We compare a number of variations with different $C_{\rm d}$ values, shown in Fig. 14.
[…]
Figure 14: Properties of dust grains in logarithmic bins of gas density - following Fig. 13, now comparing runs with varying levels of turbulent diffusion. Lines from blue to yellow indicate increasing diffusion, from no diffusion, 0.1$\times$ fiducial, fiducial to 10$\times$ fiducial. We see that stronger diffusion leads to a lower transition density from low to high $\mathcal{DTZ}$, with only a marginal change to the $S/L$ ratio. The transition density for dust (H2) decreases by 0.3 dex (0.4 dex) between the no and high diffusion coefficient cases.
@@ Appendix C Dust and Metallicity Calibrations
A curious tension in the development of these models has been the diverging conclusions on the abundances of dust in our simulated galaxies that can be drawn through comparison to different data sets. While comparisons like the cosmic dust density, galaxy dust mass functions and dust scaling relations (see Figs. 6, 7, 8) may point to a slight surfeit of dust, $\mathcal{DTG}$ and $\mathcal{DTZ}$ values seem to imply a deficit of dust in the Fiducial run compared to observation (see e.g. Fig. 9).
@@
In Fig. 16, We then plot the $\mathcal{DTG}$ values of Fiducial simulation galaxies as a function of their diffuse oxygen abundances, against the De Vis et al. (2019) data for these two calibrations. We see that the simulated data agrees much better with the $\mathcal{DTG}$ values when using the Tremonti et al. (2004)-calibrated metallicities. This demonstrates how our modelling shows a preference for certain metallicity calibrations.

### v2

@@ 4.6 Dust-to-gas ratio
[…]
@@
For $\mathcal{DTG}$, plotting values as a function of gas-phase metallicity can reveal how dust content varies with the ongoing enrichment of the ISM. Observationally, this metallicity is typically measured in the gas-phase via oxygen emission lines, taken as a proxy for total metallicity, $12+\log_{10}({\rm O/H})$. For our simulated galaxies, we compute this per subhalo for our simulated galaxies in 50 pkpc apertures about the most-bound particles, measuring the gas-phase oxygen fractions in relatively cold, dense gas ($\log_{10}T/{\rm K}<4.5$, $\log_{10}n_{\rm H}/{\rm cm^{-3}}>-1$) gas, to approximate the regions probed observationally. Here, we limit the galaxy selection to actively star-forming galaxies (with specific star formation rates ${\rm SFR}/M_{\star}>0.01\;{\rm Gyr}^{-1}$), with $12+\log_{10}({\rm O/H})$ abundances aggregated linearly over the subhalo material (such that O/H is the ratio of total oxygen to hydrogen across all the cool-dense, in-aperture gas). Another important consideration for the comparison in this case is that with $12+\log_{10}({\rm O/H})$ as the independent variable, there can be considerable blending in particle resolution across the $x$-axis, particularly due to low-mass galaxies scattered to high metallicity bins. By default, we apply a stellar mass cut of $\log_{10}(M_{\star}/{\rm M_{\odot}})>9$ (corresponding to galaxies resolved by $\gtrapprox 500$ star particles) to our primary median relation (solid line), while also plotting a median relation using a lower $\log_{10}(M_{\star}/{\rm M_{\odot}})>8$ cut ($\gtrapprox 50$ star particles) to show the influence of less well resolved systems.
@@
Regarding the scatter in the relation between $\mathcal{DTG}$ and metallicity, we first see that the intrinsic scatter in our simulated galaxies appears significantly smaller than that of the cloud of observed data points, with a $\pm 1\sigma$ range below 0.4 dex across the metallicity range for galaxies of $M_{\star}>10^{9}~{\rm M_{\odot}}$. Considering these data sets individually, we see that a lot of this global scatter comes from systematic offsets between datasets, though the majority of observed sets individually exhibit scatter that is larger than predicted [^18] The De Vis et al. (2019) dataset can be seen isolated in Appendix C.. An exception is the relation of Konstantopoulou et al. (2023), which is significantly tighter. It is possible that a lack of scatter could be due to the limited resolution of the simulations, given we are unable to directly model ISM structure below the resolution scale, that could provide an additional source of variance between galaxies. However, variation in scatter between data sets, given intrinsic uncertainties associated with metallicity calibrations and dust measurements, could contribute significantly to the larger observed scatter.
Comparing the relation for better resolved ($\gtrapprox 500$ star particles) galaxies with a broader selection using a lower resolution threshold ($\gtrapprox 50$ star particles), we see that for ($12+\log_{10}({\rm O/H})>8.4$) the broader selection is marginally lower ($\lesssim 0.1$ dex) owing to the inclusion of less-resolved galaxies. We begin to see the broad selection $\mathcal{DTG}$s diverge to lower values at $12+\log_{10}({\rm O/H})\sim 8.25$, falling super-linearly to $12+\log_{10}({\rm O/H})\lesssim 8.1$, and approximately linearly again below that. We note that this transition feature appears to be resolution-dependent, shifting to lower metallicities at higher resolution (Vijayan et al. in prep.). For the narrower selection $12+\log_{10}({\rm O/H})\lesssim 8.2$ we run out of galaxies, owing to a lack of low-metallicity massive objects.
@@ 4.7 Dust grain sizes
Another property we can probe with our dust evolution model is the grain size distribution. In Fig 11, we plot the logarithmic mass ratio of small grains relative to large grains for galaxies, $\log_{10}S/L$‘. We first focus on ratios computed for all gas within the 50 pkpc exclusive aperture defining each galaxy (blue lines, shading and points). This provides a metric for the grain size distribution in galaxies, where positive values are small-grain mass dominated and negative values are large-grain mass dominated. We see that for our Fiducial model, median galaxy dust masses are dominated by large (0.1 µm) grains for $\log_{10}(M_{\star}/{\rm M_{\odot}})>8.5$. In general, we see a weakly decreasing trend for our median values, such that the average galaxy goes from rough parity between the mass in small and large grains, to most massive galaxies having $\approx 60\%$ of their dust mass in large grains. This trend is significantly smaller than the scatter, however, which is of order 0.4-0.6 dex, and exhibiting diverse extremes; extending between close to an order of magnitude more mass in large grains, to around $50\%$ more mass in small grains. This diversity in the size distribution is profound, and will necessarily impact the extinction properties of grains (which are particularly relevant for post-processing observables from simulated galaxies), as well as the rates of dust-related cooling and heating processes at a fixed $\mathcal{DTG}$ ratio.
@@
To investigate this apparent inconsistency further, we can look at the small-to-large grain ratio within a more stringent gas selection. With $S/L$ ratios showing strong trends in gas density and phase (see e.g. Fig. 5), we over-plot medians for H2-weighted average $S/L$ values (red), as well a after imposing lower limits on the hydrogen number density, $n_{\rm H}>10\;{\rm cm}^{-3}$ and $n_{\rm H}>100\;{\rm cm}^{-3}$. We observe that considering H2-weighted gas does generally push $S/L$ to lower values, but only by $\approx 0.1$ dex for $\log_{10}(M_{\star}/{\rm M_{\odot}})>9$. Considering high-density gas selections has a stronger impact, with $n_{\rm H}>10\;{\rm cm}^{-3}$ gas 0.3-0.5 dex lower than the all-gas ratios, and $n_{\rm H}>100\;{\rm cm}^{-3}$ with an offset of >1>1 dex below all-gas, consistent with the lower sequence of observed values. Along with the gas selection, we note that again resolution appears to play a role here, with the higher m5 (lower m7) resolution yields higher (lower) $S/L$ ratios respectively (Vijayan et al. in prep.).
@@ 5 Discussion & Conclusions
We have presented a model for the life-cycle of dust and its interaction with the physics of the ISM, implemented using the SWIFT SPH code (Schaller et al., 2024), as part of a network of physics modules constituting the COLIBRE suite of galaxy formation simulations (Schaye et al., 2025). This model uses an established two-size (0.01 $\rm\upmu m$} and 0.1 $\rm\upmu m$}) grain population paradigm (Hirashita, 2015), as well as three chemical species of dust (carbonaceous grain, as well as Mg- and Fe-endmember silicate species), providing a model lightweight enough to be carried particle-by-particle in large SPH cosmological volume simulations. This dust is seeded from stellar channels, budgeting for consistency with overall metal enrichment, and can accrue mass from the gas-phase ISM (accretion), change its size distribution (shattering and coagulation) and ultimately be destroyed (sputtering, direct shock annihilation or astration). During this life-cycle, dust properties are coupled to the cooling and heating processes, accounting for the role of dust in e.g. absorbing and re-emitting ionising radiation, providing surfaces for molecule formation and depleting metals from the gas-phase that would otherwise contribute to cooling and heating processes. We presented results from (25 cMpc)3 moderate resolution ($m_{\rm gas}\approx m_{\rm DM}\approx 10^{6}\;{\rm M_{\odot}}$) simulations showing the effect of parameter and modelling choices on the distribution, depletion and evolution of dust, as well as results comparing galaxy dust distributions and scaling relations at redshift $z=0$ to observations.
Generally, we find that our model produces agreement with a variety of observed properties. Our Fiducial run exhibits a cosmic dust density ($\rho_{\rm c,dust}$) evolution in good agreement with a wide range of observations across different redshift windows and observational probes (Fig. 7). We put particular stock in the recent Chiang et al. (2025) results, which integrate the cosmic infrared background to probe all dust emission. We regard this as more representative of the $\rho_{\rm c,dust}$ value that we compute for the simulations, where we sum all the dust in the simulated volume. This shows a new level of agreement across the redshift range for cosmological models of dust evolution, compared to various literature models we consider. While it is arguable that the predicted $\rho_{\rm c,dust}$ evolution is flatter than observed over the range $0<z<2$, on the high side compared to low redshift observations and vice-versa, this is consistent with volume effects in a relatively small 253 cMpc3 box that lacks the massive halos that assemble vigorously at high redshift but are largely evacuated of ISM at low redshift.
@@ Acknowledgements
We are thankful to our referee whose detailed suggestions have helped to improve a number of parts of the manuscript significantly. We gratefully acknowledge everyone who has contributed to the COLIBRE project for their input. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure. JT acknowledges support of his STFC Early Stage Research and Development grant (ST/X004651/1). We are grateful to Maarten Baes, Andrea Gebek and Nick Andreadis for helpful discussions in the latter phase of this work. This work made generous use of the pipeline tools developed for the COLIBRE project, representing many years of work; including the swiftsimio (Borrow & Borrisov, 2020) and SOAP212121github.com/openjournals/joss-reviews/issues/7851 tools. SP acknowledges support from the Austrian Science Fund (FWF) through project V 982-N. ABL acknowledges support by the Italian Ministry for Universities (MUR) program ‘Dipartimenti Di Eccellenza 2023-2027’ within the Centro Bicocca dinCosmologia Quantitativa (BiCoQ), and support by UNIMIB’s Fondo Di Ateneo Quota Competitiva (project 2024-ATEQC-0050). CSF acknowledges support from ERC Advanced Investigator Grant, DMIDAS (GA 786910). This project has received funding from the Netherlands Organization for Scientific Research (NWO) through research programme Athena 184.034.002.
@@
Appendix A Additional Parameter Variations
A.1 Smaller small grain sizes
We see that smaller grains shift the depletion-limited $\mathcal{DTZ}$ regime to lower densities, and boost the $S/L$ grain ratio by $\approx 0.1$ dex relative to Fiducial in $2\lesssim\log_{10}n_{\rm H}/{\rm cm^{-3}}\lesssim 2$. While the change in these $\mathcal{DTZ}$-density relations is relatively small, we see that this acts in concert with the smaller grain size to produce a more marked shift in the H2 transition (vertical line mark) to lower densities; a higher number of grains with higher surface-to-volume ratios at intermediate densities speeds up H2 nucleation.
A.2 Turbulent diffusion
Another process that can influence the evolution of the dust, and its associated influence on gas physics, is turbulent diffusion. As described in 3.6, the diffusion of dust follows that of gas-phase elements described in Correa et al. (2025). The amount of diffusion is modulated by a diffusion constant, $C_{\rm d}$, where elements or grains are diffused more readily for higher values. We compare a number of variations with different $C_{\rm d}$ values, shown in Fig. 13.
At ISM densities ($n_{\rm H}\gtrsim 1\;{\rm cm^{-3}}$) we see the largest difference between the NoDiff and LoDiff runs. This shows the important role even low-level of diffusion can have in simulating dust growth; diffusion allows transport of dust into gas particles that have not been directly enriched by stellar sources. As accretion relies on the dust-gas interaction, without these low seeding levels of grains, dust cannot grow to deplete elements at the levels we observe in galaxies. This demonstrates that while the dust mass in our simulated galaxies is much greater than directly seeded (see e.g. Fig. 5), the production and distribution of seed grains is crucial to manifest the dust content of galaxies. We also note that the Hi-H2 transition density $\rho_{\,0.5,\;{\rm H_{2}}}$ differs most for the NoDiff run; with dust only residing in gas particles that have been directly enriched, the important role of dust in nucleating H molecules is lacking in neighbouring gas.
[…]
Appendix B Extinction Efficiencies and Dust Coupling
@@ Appendix C Dust and Metallicity Calibrations
A curious tension in the development of these models has been the diverging conclusions on the abundances of dust in our simulated galaxies that can be drawn through comparison to different data sets. While comparisons like the cosmic dust density, galaxy dust mass functions and dust scaling relations (see Figs. 7, 8, 9) may point to a slight surfeit of dust, $\mathcal{DTG}$ and $\mathcal{DTZ}$ values seem to imply a deficit of dust in the Fiducial run compared to observation (see e.g. Fig. 10).
@@
In Fig. 16, we then plot the $\mathcal{DTG}$ values of Fiducial simulation galaxies as a function of their gas-phase oxygen abundances, against the De Vis et al. (2019) data for these two calibrations. We see that the simulated data agrees much better with the $\mathcal{DTG}$ values when using the Tremonti et al. (2004)-calibrated metallicities. This demonstrates how our modelling shows a preference for certain metallicity calibrations.